### v1

@@ 6 Conclusion and Outlook
Looking ahead, the limitations identified in our reflections (section 5) define a clear research agenda. Future work can build on this foundation to create richer, more realistic simulations that capture the full spectrum of user behaviors, especially for conversational search. The resource provides a practical basis for testing and refining new user simulators, exploring alternative evaluation measures, and investigating how simulators generalizes across different tasks and systems. As a framework, it enables reproducible experiments, highlights the trade-offs between realism, control, and evaluative rigor, and helps the community identify best practices. By making this resource accessible to the research community, we hope to encourage broader adoption, systematic benchmarking, and iterative improvement of simulation-based methods in interactive IR. Ultimately, the experiences gained from this initiative serve as a blueprint for our long-term goal: to establish a dedicated shared track at a major venue like TREC or CLEF, focused entirely on the broader research challenges of validating user simulators.

### v2

@@ 6 Conclusion and Outlook
Looking ahead, the limitations identified in our reflections (Section 5) define a clear research agenda. Future work can build on this foundation to create richer, more realistic simulations that capture the full spectrum of user behaviors, especially for conversational search. The resource provides a practical basis for testing and refining new user simulators, exploring alternative evaluation measures, and investigating how simulators generalize across different tasks and systems. As a framework, it enables reproducible experiments, highlights the trade-offs between realism, control, and evaluative rigor, and helps the community identify best practices. By making this resource accessible to the research community, we hope to encourage broader adoption, systematic benchmarking, and iterative improvement of simulation-based methods in interactive IR. Ultimately, the experiences gained from this initiative serve as a blueprint for our long-term goal: to establish a dedicated shared track at a major venue like TREC or CLEF, focused entirely on the broader research challenges of validating user simulators.
{credits}
6.0.1 Acknowledgements
This work is partially funded by Deutsche Forschungsgemeinschaft (DFG) under grant number 509543643 and within the funding programme FH-Personal (PLan CV, reference number 03FHP109) by the German Federal Ministry of Education and Research (BMBF) and Joint Science Conference (GWK).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.